Rutner Rope – Balance Sheet

Current Assets

(Thousands)

Cash

23.2

Net Receivables

119.5

Inventories

68.6

Other Current Assets

4.6

Total Current Assets

171

Net Fixed Assets

466.8

Other Noncurrent Assets

715.7

Total Assets

1353.5

Liabilities and Shareholders' Equity

Current Liabilities

Accounts Payable

58.1

Short-Term Debt

4

Other Current Liabilities

103.8

Total Current Liabilities

155.9

Long-Term Debt

849.8

Other Noncurrent Liabilities

95.5

Total Liabilities

1320.7

Revenue

2456.3

COGS

2068.4

Gross Profit

387.9

  1. What are accounts payable in days?
  1. What are inventories in days?
  1. What are accounts receivable in days?
  1. What is Rutner’s Rope’s cash-to-cash metric?
Answer #1
accounts payable in days 10.25 days
inventories in days 12.10 days
accounts recievables in days 17.75 days
cash tot cash metric 19.6 days

accounts payable in days indicates the average number of days a company needs to pay its bills and obligations.

accounts payable in days=(accounts payable/COGS)*365

accounts payable in days=(58.1/2068.4)*365=10.25 days

inventory in days indicates the average number of days that items of inventory are held for.

inventory in days=(inventory/COGS)*365

inventory in days=(68.6/2068.4)*365=12.10 days

accounts recievables in days indicate the average length of time it takes for a company's customer to pay what they owe.

accounts recievables in days=(net recievables/sales)*365

accounts recievables in days=(119.5/2456.3)*365=17.75 days

cash to cash metric=inventory in days + accounts recievables in days - accounts payable in days

cash to cash metric=12.10+17.75-10.25=19.6 days